OpenStudy (anonymous):

A spinner is divided into 10 equal sections numbered 1 through 10. If the arrow is spun once, what is the probability it will land on a number greater than 6? A) 2/5 B) 1/2 C) 3/5 D) 3/10

OpenStudy (anonymous):

4/10

OpenStudy (allank):

probability = (number of numbers greater than 6)/10 =4/10=2/5